OpenStudy (anonymous):

Solve Please :4x + 3y = -12. Solve Please :4x + 3y = -12. @Mathematics

OpenStudy (turingtest):

sove for x or y?

OpenStudy (turingtest):

solve^^^

OpenStudy (anonymous):

Answer Key: y = -12 - 4/3 x y = -4 - 4/3 x y = -12 + 4x y = -4 + 4x

OpenStudy (anonymous):

im not sure thats y im confused

OpenStudy (turingtest):

all the equations above are "solved for y" because we made them in the form y=....whatever. the correct answer is y = -4 - 4/3 x did you need steps?

OpenStudy (anonymous):

oh !! now i see thank you i was so confused i understand now :D thanks